# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ Appendix A Proof of Directional Convergence
$\qquad-\frac{q_{\gamma}^{-1}\sqrt{n}\|Z\boldsymbol{\mu}\|}{R}-\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}q_{\gamma}^{-1}\sqrt{n}(\sqrt{n}\|\boldsymbol{\mu}\|^{2}+\|Z\boldsymbol{\mu}\|)$
$\geq\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$
$\qquad-\frac{2n\|\boldsymbol{\mu}\|^{2}}{q_{\gamma}^{2}R}\tilde{\varepsilon}-\frac{q_{\gamma}+2\tilde{\varepsilon}}{3q_{\gamma}^{2}}\sqrt{n}\varepsilon_{2},$
where the second and third inequalities are due to $\|B^{-1}\|^{2}\leq q_{\gamma}^{-1}$ and $\|q_{+}B_{+}+q_{-}B_{-}\|\leq 1$, and the last inequality is due to the definition of $\tilde{\varepsilon}_{2}$.
We split in cases to bound the first term $\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$.
Case $q_{+}=1,q_{-}=0$: In this case, the first term becomes $\|\boldsymbol{\mu}\|^{2}(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}\boldsymbol{y}$. By (36), we have
$\|\boldsymbol{\mu}\|^{2}(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}\boldsymbol{y}$ (66)
$=$ $\frac{\|\boldsymbol{\mu}\|^{2}}{R}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{+}\right.$
$\qquad\qquad\left.+\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{-}\right\}$
$\geq$ $\frac{\|\boldsymbol{\mu}\|^{2}}{R}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)(n_{+}+n_{-})$
$\geq$ $\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{(1+\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}})R}\gamma^{-1}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq$ $\left(1-\frac{\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}}{1+\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}}\gamma^{-1}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$=$ $\frac{\gamma+(1+\gamma)\tilde{\varepsilon}_{3}}{(1+\gamma)(1+\tilde{\varepsilon}_{3})}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq$ $\frac{\gamma}{2(1+\gamma)}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}.$
Thus, by (65) and noting $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$, we have for the regime (i)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle\geq\left(\frac{\gamma}{2(1+\gamma)}-\frac{2\tilde{\varepsilon}}{\gamma^{4}}-\frac{2}{3\gamma^{2}}\frac{R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\tilde{\varepsilon}_{2}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\geq\frac{\gamma}{4(1+\gamma)}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}.$
Case $q_{+}=0,q_{-}=1$: In this case, the first term becomes $\|\boldsymbol{\mu}\|^{2}(B_{-}^{-1}\boldsymbol{y})^{\top}A^{-1}\boldsymbol{y}$. Similarly, under the regime (i), (39) implies the same lower bound as the case $q_{+}=1,q_{-}=0$.
Case $q_{+}q_{-}\neq 0$: By (45) and $B^{-1}(q_{+}B_{+}\boldsymbol{y}+q_{-}B_{-}\boldsymbol{y})=\frac{q_{+}+\gamma q_{-}}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{1}_{+}-\frac{q_{-}+\gamma q_{+}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{1}_{-}$, we have
$\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$ (67)
$=$ $\|\boldsymbol{\mu}\|^{2}\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
×[$\,\times\left[\left\{\left(q_{-}+\gamma^{2}q_{+}-\gamma\right)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\right\}(q_{+}+\gamma q_{-})n_{+}\right.$
+$\,\left.+\left\{\left(q_{+}+\gamma^{2}q_{-}-\gamma\right)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\right\}(q_{-}+\gamma q_{+})n_{-}\right].$
Under the regime (ii), (48), (49), and (67) imply
$\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$ (68)
$\geq$ $\|\boldsymbol{\mu}\|^{2}\frac{\frac{q_{\gamma}R}{2}\{(q_{+}+\gamma q_{-})n_{+}+(q_{-}+\gamma q_{+})n_{-}\}}{q_{\gamma}\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}R^{2}}$
$\geq$ $\|\boldsymbol{\mu}\|^{2}\frac{\frac{q_{\gamma}^{2}R}{2}(n_{+}+n_{-})}{q_{\gamma}\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}R^{2}}$
where the second inequality follows from $\min\{q_{+}+\gamma q_{-},q_{-}+\gamma q_{-}\}\geq q_{\gamma}$
With this, (65), and $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$, we have under regime (ii)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $\geq\frac{q_{\gamma}}{2\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}-\frac{2n\|\boldsymbol{\mu}\|^{2}}{q_{\gamma}^{2}R}\tilde{\varepsilon}-\frac{q_{\gamma}+2\tilde{\varepsilon}}{3q_{\gamma}^{2}}\sqrt{n}\varepsilon_{2}$ (69)
@@
=$=\left\{\frac{q_{\gamma}}{2\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}-\frac{2}{q_{\gamma}^{2}}\tilde{\varepsilon}-\frac{q_{\gamma}+\tilde{\varepsilon}}{3q_{\gamma}^{2}}\frac{R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\tilde{\varepsilon}_{2}\right\}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
@@
$\geq\left\{\frac{q_{\gamma}}{2\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}-\frac{2}{q_{\gamma}^{2}}\tilde{\varepsilon}-\frac{2}{3q_{\gamma}}\frac{R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\tilde{\varepsilon}_{2}\right\}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
@@
$\geq\frac{q_{\gamma}}{4\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq\frac{q_{\gamma}}{4\left(1+\lambda\right)^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq\frac{q_{\gamma}}{4\left\{1+\frac{q_{\gamma}}{\gamma(1-\gamma)}\right\}^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R},$
where the last two lines are due to $\tilde{\varepsilon}\leq 1$ and $\lambda\leq\frac{q_{\gamma}}{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}$.
Under regime (iii), the definition of $\tilde{E}_{3}$, (51), and (67) imply
$\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$ (70)
$\geq$ $\|\boldsymbol{\mu}\|^{2}\left\{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\right\}^{-1}(n\|\boldsymbol{\mu}\|^{2})^{-2}$
×$\quad\times\left\{(q_{\gamma}-\gamma)(n_{+}+n_{-})\|\boldsymbol{\mu}\|^{2}+2q_{\gamma}R\right\}q_{\gamma}(n_{+}+n_{-})$
$\geq$ $\|\boldsymbol{\mu}\|^{2}\left\{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\right\}^{-1}(n\|\boldsymbol{\mu}\|^{2})^{-2}\cdot(q_{\gamma}-\gamma)q_{\gamma}n^{2}\|\boldsymbol{\mu}\|^{2}$
$=$ $\frac{q_{\gamma}(q_{\gamma}-\gamma)}{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}}=q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda).$
Thus, by (65) and noting $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$, we have under regime (iii),
@@
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $\geq q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda)-\frac{2n\|\boldsymbol{\mu}\|^{2}}{q_{\gamma}^{2}R}\tilde{\varepsilon}-\frac{2}{3q_{\gamma}}\sqrt{n}\varepsilon_{2}$ (71)
@@
$\geq\frac{q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda)}{2}$
$\geq\frac{q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda=\tfrac{\gamma}{4})}{2},$
where the last line is due to $\lambda\geq\tfrac{\gamma}{4}$.
Finally, the desired conclusion holds by noting the following inequallity
$\max\left\{\frac{n\|\boldsymbol{\mu}\|^{2}}{R},1\right\}\geq\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Lemma B.10.
@@
Then the desired conclusion follows from the basic property of sub-Gaussian distributions (see Proposition 2.5.2, Vershynin [2018]). ∎
Theorem B.11 (Detailed version of Theorem 5.1).
Under the assumptions of Lemma B.9, we have
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\exp\left(-c\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\boldsymbol{z}\|_{\psi_{2}}^{2}\{n\|\boldsymbol{\mu}\|^{2}+R\}}\right),$
where $c$ is a constant which depends only on $q_{+}$ and $\gamma$.
@@
By Lemma B.8 and B.9, we have
$\left(\frac{\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle}{\|\boldsymbol{\bar{w}}\|}\right)^{2}\geq\frac{\left(\tilde{c}_{2}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\right)^{2}}{\tilde{c}_{1}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}}=\frac{\tilde{c}_{2}^{2}}{\tilde{c}_{1}}\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Then, the desired conclusion holds from Lemma B.10. ∎
Appendix C Proof of Theorem 6.1 and 6.2
@@
$\Big{|}\Big{\langle}\frac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|},\frac{\boldsymbol{z}_{k}}{\|\boldsymbol{z}_{k}\|}\Big{\rangle}\Big{|}$ $\leq\frac{\frac{\tilde{\varepsilon}_{1}}{3}R}{(1-\frac{\tilde{\varepsilon}_{1}}{3})R}$
@@
Lemma C.2 (Lemma S4.1, S4.2, Hashimoto et al. [2025]).
@@
With these, we can rewrite assumptions of our main theorems, namely Theorem 4.4, 4.9, and B.11, using parameters defining the model (sG).
From Theorem 4.4 and 4.9, we have the following result on (sG).
Theorem C.5.
@@
Then, we have with probability at least $1-2\delta$, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given by Theorem 4.1.
@@
By Lemma C.1, C.2, and the assumption (72), all the conclusions of Lemma C.1 and C.2 hold with probability at least $1-2\delta$ with
@@
$\leq$ $1.09C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|$
$\qquad\qquad+0.6C_{1}\max\left\{\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},n\log(\tfrac{1}{\delta})\|\Sigma\|\right\}$
$\leq$ $2.18\max\{C_{1}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},C_{1}n\log(\tfrac{1}{\delta})\|\Sigma\|,C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\}.$
Under condition (i), the desired conclusions follow from Theorem 4.4.

### v2

@@ Appendix B Proof of Directional Convergence
$\|(B^{-1}\boldsymbol{y})^{\top}A^{-1}\|\leq\begin{cases}\gamma^{-1}\dfrac{\sqrt{n}}{R},&\text{if $q_{+}q_{-}=0$,}\\ \dfrac{8q_{\gamma}^{\frac{1}{2}}}{(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\dfrac{\sqrt{n}}{n\|\boldsymbol{\mu}\|^{2}+R},&\text{if $q_{+}q_{-}\neq 0$.}\end{cases}$ (93)
The remaining argument is split in cases.
Regime (i):
We first note (71) and $n\|\boldsymbol{\mu}\|^{2}\leq\gamma R$ imply
@@
$\frac{9}{64(1+\gamma)^{3}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\leq\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\leq(1+\gamma)^{2}\left(\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}.$ (94)
Furthermore, Lemma C.9 and (93) imply
@@
$\left|\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle-\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\right|$ $\leq\frac{\gamma^{-1}q_{\gamma}^{-\frac{1}{2}}}{3}\sqrt{n}\tilde{\varepsilon}_{2}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2}$ (95)
$=\frac{\gamma^{-1}q_{\gamma}^{-\frac{1}{2}}+2q_{\gamma}^{-2}\tilde{\varepsilon}}{3}\sqrt{n}\tilde{\varepsilon}_{2}$
$=\frac{\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}}{3}\sqrt{n}\tilde{\varepsilon}_{2}$
$\leq\frac{\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}}{3}\tilde{D}\frac{n\|\boldsymbol{\mu}\|^{2}}{R},$
where the last line is due to the assumption on $\tilde{\varepsilon}_{2}$.
By (94) and (95), it suffices to choose $\tilde{D}$ so that
$\frac{\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}}{3}\tilde{D}\leq\frac{1}{8(1+\gamma)^{3}}.$
Then, we have
$\frac{1}{64(1+\gamma)^{3}}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\leq\frac{1}{64(1+\gamma)^{3}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\leq\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$
@@
$\quad\leq\left\{\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}+\frac{1}{8(1+\gamma)^{3}}\right\}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\leq(1+\gamma)\left\{\gamma^{-2}+2\gamma^{-4}\tilde{\varepsilon}+\frac{1}{8(1+\gamma)^{3}}\right\}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Regime (ii): We note that under regime (ii), we have $\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\approx\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}$ by (79) and (82). On the other hand, Lemma C.9 and (93) imply
@@
$\left|\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle-\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\right|$ $\leq\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\tilde{\varepsilon}_{2}\frac{\sqrt{n}R}{n\|\boldsymbol{\mu}\|^{2}+R}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2}$ (96)
$=\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\frac{\tilde{\varepsilon}_{2}R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2}$
$\leq\frac{8\tilde{D}}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}++\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\tilde{D}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}$
$=\left\{\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\right\}\tilde{D}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Therefore, choosing an appropriate $\tilde{D}$ suffices to establish the desired bounds.
Specifically, under regime (ii), we have by (82)
$\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\geq\frac{q_{\gamma}}{5(1+\lambda)}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R},$
and hence it suffices to choose $\tilde{D}$ so that
$\left\{\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\right\}\tilde{D}\leq\frac{q_{\gamma}}{10(1+\lambda)}.$
Regime (iii): Similarly, we note that under regime (iii), we have $\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\approx\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\approx 1$ by (86) and (89). On the other hand, Lemma C.9 and (93) imply
$\left|\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle-\|\boldsymbol{\mu}\|^{2}\sqrt{K_{1}}\right|$ (97)
$\leq$ $\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\tilde{\varepsilon}_{2}\frac{\sqrt{n}R}{n\|\boldsymbol{\mu}\|^{2}+R}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2}$
$=$ $\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\frac{\tilde{\varepsilon}_{2}R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}\cdot\sqrt{n}\tilde{\varepsilon}_{2}$
$\leq$ $\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\frac{\tilde{\varepsilon}R}{n\|\boldsymbol{\mu}\|^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}^{2}$
$\leq$ $\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\lambda^{-1}\tilde{\varepsilon}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}^{2}(1+\lambda^{-1})^{-1}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}$
@@
$=$ $\left\{\frac{8}{3(1-\tilde{\varepsilon}_{3}^{2})(1-\gamma^{2})^{2}q_{+}q_{-}}\lambda^{-1}+\frac{2q_{\gamma}^{-2}}{3}\tilde{\varepsilon}(1+\lambda^{-1})^{-1}\right\}\tilde{\varepsilon}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
By (86), (89), and (97), we have the desired bound for sufficiently large $n$.
Lemma C.12.
@@
If $\boldsymbol{z}=\Sigma^{\frac{1}{2}}\boldsymbol{\xi}$ with $\boldsymbol{\xi}\sim\mathcal{N}(0,I_{p})$, then we further have
$\kappa\left(\beta_{min}^{-\frac{1}{2}}\frac{\langle\boldsymbol{w},\boldsymbol{\mu}\rangle}{\|\boldsymbol{w}\|}\right)\leq\mathbb{P}(\langle\boldsymbol{w},y\boldsymbol{x}\rangle\leq 0)=\kappa\left(\frac{\langle\boldsymbol{w},\boldsymbol{\mu}\rangle}{\|\Sigma^{\frac{1}{2}}\boldsymbol{w}\|}\right)\leq\kappa\left(\beta_{max}^{-\frac{1}{2}}\frac{\langle\boldsymbol{w},\boldsymbol{\mu}\rangle}{\|\boldsymbol{w}\|}\right),$ (98)
where $\kappa(t)=\mathbb{P}(|\xi_{1}|\geq t)$ and $\beta_{min}$ is the smallest eigenvalue of $\Sigma$ and $\beta_{max}=\|\Sigma\|$.
@@
Then the desired conclusion follows from the basic property of sub-Gaussian distributions (see Proposition 2.5.2, Vershynin [2018]).
The second result follows since $\langle\boldsymbol{w},\boldsymbol{z}\rangle=\langle\boldsymbol{w},\Sigma^{\frac{1}{2}}\boldsymbol{\xi}\rangle=\langle\Sigma^{\frac{1}{2}}\boldsymbol{w},\boldsymbol{\xi}\rangle\sim\mathcal{N}(0,\boldsymbol{w}^{\top}\Sigma\boldsymbol{w})$ if $\boldsymbol{\xi}\sim\mathcal{N}(0,\Sigma)$. ∎
Theorem C.13 (Detailed version of Theorem 5.1).
Suppose event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with one of the following conditions:
  1. (i)
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\frac{q_{\gamma}}{2\gamma(1-\gamma)}R$, $\sqrt{n}\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}}{20}$ , and $\tilde{\varepsilon}_{3}\leq\frac{1}{2}$,
  3. (iii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\frac{q_{\gamma}}{2\gamma(1-\gamma)}R$, $\tilde{\varepsilon}_{3}\leq\frac{1}{2}$,and $\sqrt{n}\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}(q_{\gamma}-\gamma)}{40}\frac{R}{n\|\boldsymbol{\mu}\|^{2}}$
Further suppose $\tilde{\varepsilon}_{2}\leq\tilde{D}\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$ in (i) and (ii), where $\tilde{D}$ is a constant which depends only on $\gamma$, $q_{+}$.
Then, there exist constants $c$ and $N$, which depend on $\gamma$ and $q_{+}$, such that $n\geq N$ imply
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\exp\left(-c\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\boldsymbol{z}\|_{\psi_{2}}^{2}\{n\|\boldsymbol{\mu}\|^{2}+R\}}\right).$
If $\boldsymbol{z}=\Sigma^{\frac{1}{2}}\boldsymbol{\xi}$ with $\boldsymbol{\xi}\sim\mathcal{N}(0,I_{p})$, then we have instead
$\kappa\left\{c^{-1}\beta_{min}^{-\frac{1}{2}}\left(\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}\right)^{\frac{1}{2}}\right\}\leq\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\kappa\left\{c\beta_{max}^{-\frac{1}{2}}\left(\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}\right)^{\frac{1}{2}}\right\},$
where $\kappa(t)=\mathbb{P}(|\xi_{1}|\geq t)$ and $\beta_{min}$ is the smallest eigenvalue of $\Sigma$ and $\beta_{max}=\|\Sigma\|$.
@@
Since we assume $\sqrt{n}\tilde{\varepsilon}\lesssim 1$ in all the regimes, we have $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$ for sufficiently large $n$.
By Lemma C.10 and C.11 with $\lambda=\frac{q_{\gamma}}{2\gamma(1-\gamma)}$, we have
$\frac{\hat{c}_{2}^{2}}{\tilde{c}_{1}}\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}\leq\left(\frac{\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle}{\|\boldsymbol{\bar{w}}\|}\right)^{2}\leq\frac{\tilde{c}_{2}^{2}}{\hat{c}_{1}}\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Then, the desired conclusion holds from Lemma C.12. ∎
Appendix D Proof of Theorem 6.2 and Theorem 6.3
@@
[…]
@@
Lemma D.2 (Lemma S4.1, S4.2, Hashimoto et al. [2025]).
@@
With these, we can rewrite assumptions of our main theorems, namely Theorem 4.8 and C.13, using parameters defining the model (sG).
From Theorem 4.8, we have the following result on (sG).
Theorem D.5.
@@
Then, we have with probability at least $1-2\delta$, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given in Theorem 4.8.
@@
By Lemma D.1, D.2, and the assumption (99), all the conclusions of Lemma D.1 and D.2 hold with probability at least $1-2\delta$ with
@@
[…]
We first show the desired conclusions under condition (i).